OpenStudy (firejay5):

Find the surface area of each Prism. Show work and explain. Medal will be rewarded for patronage! :D

OpenStudy (firejay5):

Instead of 10 cm, it should be 4 in and instead of 25 cm = 8 in

OpenStudy (firejay5):

@surjithayer @whpalmer4

OpenStudy (anonymous):



OpenStudy (firejay5):

I said change the 10 to 4 inches and change the 25 to 8 inches

OpenStudy (anonymous):

SIMILARLY YOU CAN DO. 

OpenStudy (firejay5):

? = 8

OpenStudy (firejay5):

I meant 80

OpenStudy (anonymous):

\[=32\sqrt{2}+64+16=80+32\sqrt{2}\]